Using the data in Appendix 2 of your bestbook. Cal
Answer #1

For the reaction CuO(s)+ H2(g)--àCu(s)+ H2O(g)

Entropy change = sum of entropy of products –sum of entropy of reactants

If S denotes entropy

Entropy change = 1*SCu +1*S(H2O(g) – {1*S(cuO) +1*S(H2) }

Where 1 are the coefficients of all the reactants and products

Entropy data : S(Cu(s))= 33.3 J/mol.K, S (H2O(g) = 188.83 J/mol.K, S(CuO (s))= 42.59 J/mol.K and S(H2(g)= 130.58 J/mol.K)

Standard Entropy change = 33.3+188.83- (42.59+130.58)= 49.96 J/mol.K

2. for the reaction, 2Al+3H2O(g) --à Al2O3(s)+ 3H2(g)

S(Al2O3(s)= 50.9 J/mol.K, S(Al)= 28.3 J/mol.K, S(H2O(g)= 188.83 J/mol.K, S(H2(g)= 130.58 J/mol.K

Entropy change = S(Al2O3)+3* S(H2)- {2*S(Al)+3*S(H2O(g)}

1,3, 2 and 3 are coefficients of respective components

Standard Entropy change = 50.9+3*130.58-(2*28.3+3*188.83)= -179.55 J/K

3. CH4(g) +2O2(g)--à CO2(g)+2H2O(l)

Standard entropy change = S(CO2(g)+2*S(H2O(l) –{ S(CH4,(g)+ 2*S(O2(g)}

1,2, 1 and 2 are coefficients of components in the reactions.

S(O2)= 205 J/mol.K, S(CH4)= 186.2 J/mol. K, S(CO2)= 213.6, S(H2O,(l)= 69.9 J/mol.K

=213.6+2*69.9- {186.2+2*205} =-242.8 J/K